OpenStudy (anonymous):

What is the slope of the coordinates (1,3) and (2,-1)

OpenStudy (jim766):

do you know the slope formula?

OpenStudy (anonymous):

m=y1-y2/x2-x1

OpenStudy (anonymous):

(x,y)

OpenStudy (anonymous):

good so far?

OpenStudy (anonymous):

ya but i dont get how to plug it in

OpenStudy (anonymous):

your (x,y) are your numbers so just assign x1 and y1 to first bracket and x2 and y2 to second

OpenStudy (anonymous):

(x1,y1)= (1,3) (x2,y2)=?

OpenStudy (anonymous):

so is it 3-(-1)/1-2 ?

OpenStudy (anonymous):

So if thats the slope then what would be the equation?

OpenStudy (jim766):

first lets finish the slope the numerator would be 4 what would the denominator be?

OpenStudy (anonymous):

-1

OpenStudy (jim766):

right so the slope or m = 4/-1 or -4

OpenStudy (anonymous):

ya i got that but what do i do to get the equation because i already have the slope

OpenStudy (jim766):

In order to make a linear equation y = mx + b, you need a slope and a y intercept. We have a slope now, but need a y intercept. Pick one of the points, it doesnt matter which one.

OpenStudy (jim766):

then put the x and the y value into y = (-4)x + b Then solve for b

OpenStudy (jim766):

Which point do you want to use...like I said it doesnt matter which one.

OpenStudy (anonymous):

(1,3)

OpenStudy (jim766):

(x, y ) (1, 3) now substitute the x and the y into y = (-4)x + b

OpenStudy (anonymous):

b = -1?

OpenStudy (jim766):

y = (-4)x + b 3 = (-4)(1) + b 3 = -4 + b +4 +4 7 = b what did you do wrong?

OpenStudy (anonymous):

when i moved the 4

OpenStudy (anonymous):

i forgot it cancels that four out so i got it mixed

OpenStudy (jim766):

right, always make sure you are doing the opposite so now we know m = -4 b = 7 all you have to do is put them into y = mx + b y = -4x + 7 do you see?

OpenStudy (anonymous):

yes

OpenStudy (jim766):

when you are finding an equation this way and you have two points, pick the one with what look like the easiest numbers, in this case both points were similar. Have a good one!

OpenStudy (anonymous):

thanks